OpenStudy (anonymous):

Hi I really need help with a math problem. Could somebody PLEASE help me solve this??!?!??!

OpenStudy (anonymous):

what is it?

OpenStudy (anonymous):

{y-3x=6 20x^2-22x-y=-11 Solve the following systmes of equations^^^^ PLEASE HELP

OpenStudy (anonymous):

could you help?

OpenStudy (anonymous):

here again substitute y=6+3x in the second equation, we get, 20x^2-25x=-5 which is 20x^2 -25x+5...my friend lagrange will factorise it for u(;) :D) from here. find the value of x and coresponding values for y.

OpenStudy (anonymous):

Thank you so much I cant wait to see what langarange comes up with because if I pass this homework I pass the class!!! Thank you so much!

OpenStudy (anonymous):

ehi! understand these...they form the basics for ur higher clases :)

OpenStudy (anonymous):

I am trying so hard but these problems are just tripping me up Could you factor it for me since it looks like your friend isnt?

OpenStudy (anonymous):

5*(4x^2-5x+1)=0 5*(4x-1)(x-1)=0

OpenStudy (anonymous):

x=1,1/4, find the corresponding values of y from the first equation

OpenStudy (anonymous):

How would you do that??

OpenStudy (anonymous):

well x=1 is a solution, so plug into y-3x=6 y-3=6 y=9 for x=1/4: y-3/4=6 y=27/4 so your solutions are (1,9) and (1/4, 27/4)

OpenStudy (anonymous):

Your amazing WOW

OpenStudy (anonymous):

let me check the answers one sec

OpenStudy (anonymous):

Thank you so much Ive got three more problems like it.. I will try these and see if i can get the answer

OpenStudy (anonymous):

It says there correct, I use an online hw thing

OpenStudy (anonymous):

yea my friend rsvitale helped u , thanks man ;) :D

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

oh allright then i wont check them haha

OpenStudy (anonymous):

Ok yet again Im lost... UGH